OpenStudy (anonymous):

The demand for a certain brand of portable alarm function is given by: p = - 0,01.x^2 - 0,3x + 10 Where p is the unit price at the wholesale real ex, the quantity demanded each month, measured in units of thousands. The bid for this watch brand function is given by: p = - 0,01x^2 + 0,2x + 4 Having po same meaning as before and where x amount, in thousands, that the supplier will market monthly. Determine the consumer surplus and the surplus production when the unit market price is equal to the equilibrium price.

OpenStudy (amistre64):

what defines equilibrium price?

OpenStudy (amistre64):

i believe equi price is when demand is equal to supply - 0,01.x^2 - 0,3x + 10 = - 0,01x^2 + 0,2x + 4 - 0,3x + 10 = 0,2x + 4 10 = 0,5x + 4 6 = 0,5x 12 = x

OpenStudy (amistre64):

but i cant be sure that my definintion for equi price is valid

OpenStudy (anonymous):

understood because they speak the demand function is given by p = - 0,01.x^2 - 0,3x + 10 and the supply function is given by: p = - 0,01x^2 + 0,2x + 4 dai ele quer saber o excedente do consumo e o excedente de produção, quando o preço unitário de mercado é igual ao preço de equilíbrio, fiquei meio confuso em relação a esse exercício...